OpenStudy (anonymous):

87x + 93y = 1092 and 87x + 93y = 12 Solve for x. Then solve for y.

OpenStudy (anonymous):

So, what do you think your first step should be?

OpenStudy (anonymous):

ok for the problem on the left, you have to cancel out to each side. 87x cancels 87x then 1092 divided by 87 = 12.5

OpenStudy (anonymous):

12.5 would be x.. no?

OpenStudy (anonymous):

sorry, hold on one second

OpenStudy (anonymous):

If you are trying to solve for both x and y given these two equations, then there is no answer. 87x + 93y cant equal both 1092 and 12

OpenStudy (anonymous):

yeah, i was gonna suggest a typo

OpenStudy (anonymous):

so r u trying to solve for x in the equation the left

OpenStudy (anonymous):

or for both x and y

OpenStudy (anonymous):

ok let me give you the word problem. 1 sec

OpenStudy (anonymous):

The octane rating of a gasoline is a measure of the amount of isooctane in the gas. Subaru recommends 91-octane gasoline for the 2008 Legacy 3.0 R. How much 87-octane gas and 93-octane gas should Kelsey mix in order to make 12 gal of 91-octane gas for her Legacy?

OpenStudy (anonymous):

the teacher said this afterwords....... "On your last problem, your equation of 87x + 93y = 1092 is correct, but if you have 2 variables, you must have 2 equations. I only see one equation. The other equation should be the sum = 12."

OpenStudy (anonymous):

Right

OpenStudy (anonymous):

So, the X and Y in this equation represent the number of gallons of each 87-octane and 93-octane fuel.

OpenStudy (anonymous):

So, we only want 12 gallons overall, no matter how much we mix with what; in that case, X + Y = 12.

OpenStudy (anonymous):

But we also want the average of our new fuels to be 91 octane: 87 * X + 93 * Y = 12 * 91.

OpenStudy (anonymous):

So, we've got the two equations to solve: X + Y = 12 87 * X + 93 * Y = 1092

OpenStudy (anonymous):

Now, where would we go from here?

OpenStudy (anonymous):

right the two equations is what she was saying.

OpenStudy (anonymous):

you have to get x by itself correct?

OpenStudy (anonymous):

Right. We also have a choice of which equation to start out from by making X by itself; which one would you choose?

OpenStudy (anonymous):

x first then y

OpenStudy (anonymous):

OK, good. So how would you get X by itself on the first equation?

OpenStudy (anonymous):

take 87x over 87 and x is by itself

OpenStudy (anonymous):

then you have 1092 divided by 87

OpenStudy (anonymous):

Right - in that case, you'd need to remember to divide Y by the same amount and move it over to the other side, too.

OpenStudy (anonymous):

but when you do that you are left with 93y = 12.5x

OpenStudy (anonymous):

divide 93 by 87 also?

OpenStudy (anonymous):

But there's actually a much easier way to do it; what happens if you solve for X in the other equation, X + Y = 12?

OpenStudy (anonymous):

easier way is good becuase i have tried this problem about 5 different times and i come up with a different answer every time

OpenStudy (anonymous):

*because

OpenStudy (anonymous):

lol

OpenStudy (anonymous):

Half of math is finding the hidden shortcuts in problems.

OpenStudy (anonymous):

So, how would you solve for X in X + Y = 12?

OpenStudy (anonymous):

good question

OpenStudy (anonymous):

Well, how would you get X by itself on one side?

OpenStudy (anonymous):

take out y,

OpenStudy (anonymous):

y divided by y cancels out then 12 divided by y

OpenStudy (anonymous):

?

OpenStudy (anonymous):

Well, are we trying to divide Y or subtract Y?

OpenStudy (anonymous):

i thought you would divide. to equal it out. I c im wrong

OpenStudy (anonymous):

If we were multiplying X and Y together to get 12, we'd definitely want to divide it out.

OpenStudy (anonymous):

oooo ooooo oooo

OpenStudy (anonymous):

so we are going to subtract now? first?

OpenStudy (anonymous):

But since we're ADDING Y to X, we want to get rid of Y by doing the opposite of adding - subtracting.

OpenStudy (anonymous):

Yep! You got it.

OpenStudy (anonymous):

So what does the equation look like now after you did that?

OpenStudy (anonymous):

ok..... 87x + 93y = 1092 -87x -87x ---------------------- 93y = 1005x

OpenStudy (anonymous):

close or did I really mess myself up?

OpenStudy (anonymous):

Close! One big thing to remember in algebra is that different letters don't like each other.

OpenStudy (anonymous):

So, we'd have to leave 1092 by itself - in this case, we'd leave it as 1092 - 87x.

OpenStudy (anonymous):

and you get 1005

OpenStudy (anonymous):

Well, the thing is that 87x means 87 multiplied by some number we don't know.

OpenStudy (anonymous):

correct

OpenStudy (anonymous):

X could be zero, X could be a billion - so unless we know what X is, we don't know what any number subtracted by it is, either

OpenStudy (anonymous):

right

OpenStudy (anonymous):

But we know that 1092 - 87 * X is exactly that - we want to leave it that way for now.

OpenStudy (anonymous):

ok

OpenStudy (anonymous):

So, if X was 1, then 1092 - 87 X would equal 1005

OpenStudy (anonymous):

But if X was 0, then 1092 - 87X would equal 1092

OpenStudy (anonymous):

ok

OpenStudy (anonymous):

so x is 1

OpenStudy (anonymous):

Del, I was reading this, and you're really good! You're like a legit teacher, the good kind who are patient and really care if you 'get it' or not. I applaud you and bedeck you with medals. :)

OpenStudy (anonymous):

Aww, thanks :p.

OpenStudy (anonymous):

yahoo go del

OpenStudy (anonymous):

ok now the problem at hand

OpenStudy (anonymous):

Right.

OpenStudy (anonymous):

I think it might be a good idea to step back a bit. Remember when I said there's sometimes an easy way to figure things out in math?

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

Let's go back to the first equation, X + Y = 12.

OpenStudy (anonymous):

How would we get X by itself, just in this single and only equation?

OpenStudy (anonymous):

subtract

OpenStudy (anonymous):

Sorry, I just got lagged out.

OpenStudy (anonymous):

ok i am lost

OpenStudy (anonymous):

Right, so what would the new equation be?

OpenStudy (anonymous):

So, we need to subtract Y from both sides of the equation X + Y = 12.

OpenStudy (anonymous):

x + y = 12 -y -y ------------- x = 11

OpenStudy (anonymous):

x= 12

OpenStudy (anonymous):

Close! Remember, letters and numbers don't like getting together willy-nilly.

OpenStudy (anonymous):

Remember when I told you that you couldn't take 87X from 1092?

OpenStudy (anonymous):

It's like that. Let's see what happens when we keep X, Y, and the numbers apart:

OpenStudy (anonymous):

X + Y = 12 - Y -Y -------------- X = 12 - Y

OpenStudy (anonymous):

Now, pick any two numbers that add up to twelve.

OpenStudy (anonymous):

6,6

OpenStudy (anonymous):

OK, good. So, just for the heck of it, let X = 6 and Y = 6.

OpenStudy (anonymous):

Then X + Y = 12, right?

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

Right! So let's look at the new equation we just found out, X = 12 - Y. Then 6 = 12 - 6. It makes sense!

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

Take another pair of numbers, 4 and 8. Since 4 + 8 = 12, then 4 = 12 - 8.

OpenStudy (anonymous):

In a way, then, X + Y = 12 and X = 12 - Y are saying the same exact thing, just with the numbers shifted around.

OpenStudy (anonymous):

ok

OpenStudy (anonymous):

That's why they're equivalent.

OpenStudy (anonymous):

Do you think you get why X = 12 - Y in the problem?

OpenStudy (anonymous):

yes because y has to be away from x

OpenStudy (anonymous):

Right - and we take Y away from X by subtracting it, because Y was added to X.

OpenStudy (anonymous):

Congrats! We're done Part 1 of the problem! (2 more parts to go).

OpenStudy (anonymous):

I'm sorry if I'm going a bit slow and not throwing the answer at you, I just want to make sure you get why you're doing stuff and not just doing it.

OpenStudy (anonymous):

well whats confusing me is im used to see it layed out as i was typing it.. not in word form